OpenStudy (anonymous):

if 3!=3*2*1, calculate the score from 1*1! +2*2! +3*3! +...+9*9!

OpenStudy (anonymous):

hint\[n\times n!=(n+1-1)\times n!=(n+1)!-n!\]

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

using mukushla hint : 1*1! = (1+1)! - 1! 2*2! = (2+1)! - 2! so far : (1+1)! - 1! + (2+1)! - 2! 2! - 1! + 3! - 2! = -1! + 3! do you see the pattern ?